U.S. Steel is considering a plant expansion to produce austenitic, precipitation hardened, duplex, and martensitic stainless steel round bars that is expected to cost $17 million now and another $10 million 1 year from now. If total operating costs will be $1.5 million per year starting 1 year from now, and the estimated salvage value of the plant is virtually zero, how much must the company make annually in years 1 through 11 to recover its investment plus a return of 17% per year?

The company must make $_______ million annually in years 1 through 11 to recover its investment plus a return of 17% per year.
Answer #1

Present Worth (PW) of costs ($ Million) = 17 + 10 x P/F(17%, 1) + 1.5 x P/A(17%, 11) = 10 + 10 x 0.8547** + 1.5 x 4.8364**

= 10 + 8.547 + 7.255

= 25.802

Annual revenue ($ million) = PW of costs / P/A(17%, 11) = 25.802 / 4.8364** = 5.33

**From P/F and P/A Factor tables